# Statistical Taylor Expansion111

### v1

@@ 2.4 Distributional Zero and Distributional Pole
Let $\tilde{y}=f(\tilde{x})$ be a strictly monotonic function, so that $\tilde{x}=f^{-1}(\tilde{y})$ exists. Formula (2.10) is the uncertainty distribution density function of $\tilde{y}$ [2] [5]. In Formula (2.10), the same distribution can be expressed in either $\tilde{x}$ or $\tilde{y}$ or $\tilde{z}$ , which are different representations of the same underlying random variable. Using Formula (2.10), Formula (2.11) gives the $\rho(\tilde{y},y,\delta y)$ for $x^{c}$.
$N(\tilde{z})d\tilde{z}=\rho(\tilde{x},x,\delta x)d\tilde{x}$ $=\rho(f^{-1}(\tilde{y}),x,\delta x)\frac{d\tilde{x}}{d\tilde{y}}d\tilde{y}=\rho(\tilde{y},y,\delta y)d\tilde{y};$ (2.10)
@@
Viewed in the $f^{-1}(\tilde{y})$ coordinate, $\rho(\tilde{y},y,\delta y)$ is Gaussian modulated by $\frac{d\tilde{x}}{d\tilde{y}}=1/f^{(1)}_{x}$, in which $f^{(1)}_{x}$ is the first derivative of $f(x)$ to $x$. A distributional zero of the uncertainty distribution happens when $f^{(1)}_{x}=\infty\rightarrow\rho(\tilde{y},y,\delta y)=0$, while a distributional pole happens when $f^{(1)}_{x}=0\rightarrow\rho(\tilde{y},y,\delta y)=\infty$. Zeros and poles have different impacts on the properties of of the uncertainty distributions.
  • If $y=f(x)$ is a linear transform of $x$, $f^{(1)}_{x}$ is a constant, and $\rho(\tilde{y},y,\delta y)$ is known to be Gaussian [5]. From another perspective, a linear transformation generates neither distributional zero nor distributional pole according to Formula (2.10).
  • Figure 2.4 shows the probability density function for $(x\pm 1)^{2}$ according to Formula (2.11), which has a distributional pole at $x=0$ that separates two defined region $x\geq 0$ and $x\leq 0$. $(0\pm 1)^{2}$ is the $\chi^{2}$ distribution [2]. At the distributional pole, the probability density function resembles a Delta distribution.
@@
In both Figure 2.4 and 2.4, the probability density function in the $\tilde{z}$ representation becomes more Normal-like when the mode of the distributions is further away from either distributional pole or distributional zero.
Formula (2.10) gives the uncertainty distribution of an analytic function. However, normal scientific and engineering calculations usually do not care about the result distribution, but just few statistics of the result, such as the result mean and deviation [2] [3]. Such simple statistics can be obtained using Taylor expansion.
@@ 2.5 Statistical Taylor Expansion
Formula (2.12) defines variance momentum $\eta(n)$, which has a simple form when the uncertainty distribution $\rho(\tilde{x},x,\delta x)$ is Gaussian. Formula (2.13) holds when the uncertainty distribution is symmetric.
An analytic function $f(x)$ can be precisely calculated in a range by the Taylor series as shown in Formula (2.14). Using Formula (2.7), (2.12), and (2.13), Formula (2.15) and (2.16) [1] gives the mean $\overline{f(x)}$ and the variance $\delta^{2}f(x)$, respectively. $\overline{f(x)}-f(x)$ is defined as the uncertainty bias which is the effect of the input uncertainty to the result.
@@
$\delta^{2}f(x)=$ $\overline{(f(x)-\overline{f(x)})^{2}}=\overline{f(x)^{2}}-\overline{f(x)}^{2}$ (2.16)
$=\sum_{n=0}^{\infty}(\delta x)^{2n}\left(\zeta(2n)\sum_{j=1}^{2n-1}\frac{f^{(j)}_{x}}{j!}\frac{f^{(2n-j)}_{x}}{(2n-j)!}-\sum_{j=1}^{n-1}\frac{f^{(2j)}_{x}\zeta(2j)}{(2j)!}\frac{f^{(2n-2j)}_{x}\zeta(2n-2j)}{(2n-2j)!}\right);$
Formula (2.15) is deduced from Formula (2.14) by $\tilde{x}^{2n}\rightarrow\zeta(2n)(\delta x)^{2n}$ and $\tilde{x}^{2n+1}\rightarrow 0$. Formula (2.16) is similarly deducted from the square of Formula (2.14).
The result variance in statistical Taylor expansion show the nature of the calculation, such as $\delta x\rightarrow P(e^{x})$, $\delta x\rightarrow\delta\sin(x)$, $P(x)\rightarrow P(x^{c})$, and $P(x)\rightarrow\delta\ln(x)$. To keep the result precision finer than $1$ for $e^{x\pm\delta x}$, $\delta x\lesssim 0.833$ regardless of $x$. In contrast, Figure 3.4 will show that the result upper bound uncertainty $\overline{\log(x\pm\delta x)}$ is $0.28\pm 0.026$. Both are expected from the relation $\delta x\rightarrow P(e^{x})$, and $P(x)\rightarrow\delta\ln(x)$.
Due to uncorrelated uncertainty assumption, Formula (2.18) and (2.19) calculate the mean and variance of Taylor expansion in Formula (2.17):
@@
$=$ $\sum_{m=0}^{\infty}\sum_{n=0}^{\infty}(\delta x)^{2m}(\delta y)^{2n}\frac{\zeta(2m)\zeta(2n)f^{(2m,2n)}_{(x,y)}}{(2m)!(2n)!};$
$\delta^{2}f(x,y)=$ $\overline{f(x,y)^{2}}-\overline{f(x,y)}^{2}=\sum_{m=0}^{\infty}\sum_{n=0}^{\infty}(\delta x)^{2m}(\delta y)^{2n}$ (2.19)
$\left(\zeta(2m)\zeta(2n)\sum_{i=0}^{2m}\sum_{j=0}^{2n}\frac{f^{(i,j)}_{(x,y)}}{i!\;j!}\frac{f^{(2m-i,2n-j)}_{(x,y)}}{(2m-i)!\;(2n-j)!}\right.$
$\;\;\;\;\;\;\left.-\sum_{i=0}^{m}\sum_{j=0}^{n}\frac{\zeta(2i)\zeta(2j)f^{(2i,2j)}_{(x,y)}}{(2i)!(2j)!}\frac{\zeta(2m-2i)\zeta(2n-2j)f^{(2m-2i,2n-2j)}_{(x,y)}}{(2m-2i)!(2n-2j)!}\right);$
Although Formula (2.19) is only for 2-dimensional, it can be extended to any dimensional.
[…]
@@ 2.6 One Dimensional Examples
$\frac{\overline{e^{x}}}{e^{x}}$ $=1+\sum_{n=1}^{\infty}(\delta x)^{2n}\zeta(2n)\frac{1}{(2n)!};$ (2.21)
$\frac{\delta^{2}e^{x}}{(e^{x})^{2}}$ $=\sum_{n=1}^{\infty}(\delta x)^{2n}\left(\sum_{j=1}^{2n-1}\frac{\zeta(2n)}{j!\;(2n-j)!}-\sum_{j=1}^{n-1}\frac{\zeta(2j)}{(2j)!}\frac{\zeta(2n-2j)}{(2n-2j)!}\right);$ (2.22)
Formula (2.24) and (2.25) give the mean and variance for $\ln(x)$, respectively:
$\ln(x+\tilde{x})-\ln(x)=\ln(1+\frac{\tilde{x}}{x})=\sum_{j=1}^{\infty}\frac{(-1)^{j+1}}{j}\frac{\tilde{x}^{j}}{x^{j}};$ (2.23)
$\overline{\ln(x)}$ $=\ln(x)-\sum_{n=1}^{+\infty}P(x)^{2n}\frac{\zeta(2n)}{2n};$ (2.24)
$\delta^{2}\ln(x)$ $=\sum_{n=1}^{+\infty}P(x)^{2n}\left(\sum_{j=1}^{2n-1}\frac{\zeta(2n)}{j(2n-j)}-\sum_{j=1}^{n-1}\frac{\zeta(2j)}{2j}\frac{\zeta(2n-2j)}{2n-2j}\right);$ (2.25)
$\sin(x+\tilde{x})=\sum_{n=0}^{\infty}\sin(x)(-1)^{n}\frac{\tilde{x}^{2n}}{(2n)!}+\sum_{n=0}^{\infty}\cos(x)(-1)^{n}\frac{\tilde{x}^{2n+1}}{(2n+1)!};$ (2.26)
$\overline{\sin(x)}$ $=\sin(x)+\sin(x)\sum_{n=0}^{\infty}(\delta x)^{2n}(-1)^{n}\frac{\zeta(2n)}{(2n)!};$ (2.27)
$\delta^{2}\sin(x)$ $=\sum_{n=1}^{\infty}(\delta x)^{2n}(-1)^{n-1}$
$\left(\cos(x)^{2}\sum_{j=0}^{n-1}\frac{\zeta(2n)}{(2j+1)!(2n-2j-1)!}-\sin(x)^{2}\sum_{j=1}^{n-1}\frac{\zeta(2n)-\zeta(2j)\zeta(2n-2j)}{(2j)!(2n-2j)!}\right);$ (2.28)
@@
$\frac{\overline{x^{c}}}{x^{c}}$ $=1+1\sum_{n=1}^{\infty}P(x)^{2n}\zeta(2n)\begin{pmatrix}c\\ 2n\end{pmatrix};$ (2.30)
$\frac{\delta^{2}x^{c}}{(x^{c})^{2}}$ $=\sum_{n=1}^{\infty}P(x)^{2n}\left(\sum_{j=1}^{2n-1}\zeta(2n)\begin{pmatrix}c\\ j\end{pmatrix}\begin{pmatrix}c\\ 2n-j\end{pmatrix}-\sum_{j=1}^{n-1}\zeta(2j)\begin{pmatrix}c\\ 2j\end{pmatrix}\zeta(2n-2j)\begin{pmatrix}c\\ 2n-2j\end{pmatrix}\right);$ (2.31)
@@ 2.7 Dependency Tracing
  • Formula (2.32) shows $\delta^{2}(f+g)$, whose dependence tracing can be demonstrated by $\delta^{2}(f-f)=0$, and $\delta^{2}(f(x)+g(y))=\delta^{2}f+\delta^{2}g$, with the latter corresponding to Formula (2.5). Formula (2.36) shows that Formula (2.16) uses Formula (2.32) between any two terms in the Taylor expansion of Formula (2.14). In contrast, Formula (2.37) shows that statistical Taylor expansion can no longer be treated as polynomial.
@@ 2.8 Traditional Execution and Dependency Problem
$\overline{f(g(x))}-f(g(x))=$ $o((\delta x)^{4})+(\delta x)^{2}\frac{1}{2}\left(f^{(2)}_{x}(g^{(1)}_{x})^{2}+f^{(1)}_{x}g^{(2)}_{x}\right);$ (2.38)
$\overline{f(g)|_{g(x)}}-f(g(x))=$ $o((\delta g)^{4})+(\delta x)^{2}\frac{1}{2}f^{(2)}_{x}(g^{(1)}_{x})^{2};$ (2.39)
$\delta^{2}f(g(x))=$ $o((\delta x)^{6})+(\delta x)^{2}(f^{(1)}_{x}g^{(1)}_{x})^{2}+$ (2.40)
$(\delta x)^{4}\left(f^{(1)}_{x}f^{(3)}_{x}(g^{(1)}_{x})^{4}+4f^{(1)}_{x}f^{(2)}_{x}(g^{(1)}_{x})^{2}g^{(2)}_{x}+(f^{(1)}_{x})^{2}g^{(1)}_{x}g^{(3)}_{x}+\right.$
$\;\;\;\;\;\;\;\;\;\;\;\;\left.\frac{1}{2}(f^{(2)}_{x})^{2}(g^{(1)}_{x})^{4}+\frac{1}{2}(f^{(1)}_{x})^{2}(g^{(2)}_{x})^{2}\right);$
$\delta^{2}f(g))|_{g(x)}=$ $o((\delta x)^{6})+(\delta x)^{2}(f^{(1)}_{x}g^{(1)}_{x})^{2}+$ (2.41)
$(\delta x)^{4}\left(f^{(1)}_{x}f^{(3)}_{x}(g^{(1)}_{x})^{2}+\frac{1}{2}(f^{(2)}_{x})^{2}(g^{(1)}_{x})^{2}\right);$
$\overline{(\sqrt{x})^{2}}-x$ $\simeq o((\delta x)^{4})+\frac{1}{4}(\delta x)^{2};$ (2.42)
$\overline{\sqrt{x^{2}}}-x$ $\simeq o((\delta x)^{4})-\frac{1}{2}(\delta x)^{2};$ (2.43)
$\delta^{2}y^{2}|_{y=\sqrt{x}}$ $\simeq o((\delta x)^{6})+(\delta x)^{2}+4(\delta x)^{4};$ (2.44)
$\delta^{2}\sqrt{y}|_{y=x^{2}}$ $\simeq o((\delta x)^{6})+(\delta x)^{2}+(\delta x)^{4};$ (2.45)
@@
  • Traditionally, an analytic expression can be broken down into simpler and independent arithmetic operations of negation, addition, multiplication, division, and etc  [^6] Sometimes square root is also regarded as an arithmetic operation.. But such approach causes the dependency problem both in both floating-point arithmetic and interval arithmetic. It also causes the dependency problem in statistical Taylor expansion. For example, if $x^{2}-x$ is calculated as $x^{2}-x$, $x(x-1)$, and $(x-\frac{1}{2})^{2}-\frac{1}{4}$, only $(x-\frac{1}{2})^{2}-\frac{1}{4}$ gives the correct result, while the other two give wrong results for wrong independence assumptions between $x^{2}$ and $x$, or between $x-1$ and $x$, respectively. Only statistical Taylor expansion always give correct results.
  • Traditionally, a large calculation may be broken into sequential steps, such as calculating $f(g(x))$ as $f(y)|_{y=g(x)}$. But this method generally no longer works in statistical Taylor expansion due to dependency tracing of $g(x)$ inside $f(g(x))$. According to Formula (2.38), (2.39), (2.40), and (2.41), in statistical Taylor expansion, $f(g(x))\neq f(y)|_{y=g(x)}$, with the latter suffers from the dependency problem according to Formula (2.42), (2.43), (2.44) and (2.45).

### v2

@@ 2.2 Distributional Zero and Distributional Pole
Let $\rho(\tilde{x},x,\delta x)$ be the distribution density function of random variable $\tilde{x}$ with mean $x$ and deviation $\delta x$. Let $\rho(\tilde{z})$ be the normalized form of $\rho(\tilde{x},x,\delta x)$, in which $\tilde{z}\equiv\frac{\tilde{x}-x}{\delta x}$, e.g., Normal distribution $N(\tilde{z})$ is the normalized form of Gaussian distribution.
Let $\tilde{y}=f(\tilde{x})$ be a strictly monotonic function, so that $\tilde{x}=f^{-1}(\tilde{y})$ exists. Formula (2.5) is the distribution density function of $\tilde{y}$ [2][5]. In Formula (2.5), the same distribution can be expressed in either $\tilde{x}$ or $\tilde{y}$ or $\tilde{z}$ , which are different representations of the same underlying random variable. Using Formula (2.5), Formula (2.6) gives the $\rho(\tilde{y},y,\delta y)$ for $x^{c}$ if $\rho(\tilde{x},x,\delta x)$ is Gaussian.
$\rho(\tilde{z})d\tilde{z}=$ $\rho(\tilde{x},x,\delta x)d\tilde{x}=\rho(f^{-1}(\tilde{y}),x,\delta x)\frac{d\tilde{x}}{d\tilde{y}}d\tilde{y}=\rho(\tilde{y},y,\delta y)d\tilde{y};$ (2.5)
@@
Viewed in the $f^{-1}(\tilde{y})$ coordinate, $\rho(\tilde{y},y,\delta y)$ is $\rho(\tilde{x},x,\delta x)$ modulated by $\frac{d\tilde{x}}{d\tilde{y}}=1/f^{(1)}_{x}$, in which $f^{(1)}_{x}$ is the first derivative of $f(x)$ to $x$. A distributional zero of the uncertainty distribution happens when $f^{(1)}_{x}=\infty\rightarrow\rho(\tilde{y},y,\delta y)=0$, while a distributional pole happens when $f^{(1)}_{x}=0\rightarrow\rho(\tilde{y},y,\delta y)=\infty$. Zeros and poles provides the strongest local modulation to $\rho(\tilde{x},x,\delta x)$:
  • If $\tilde{y}=\alpha+\beta\tilde{x}$, the result distribution is same as the original distribution, as $\rho(\tilde{y},y,\delta y)=\rho(\tilde{y},\alpha+\beta x,\beta\delta x)$ [5]. A linear transformation generates neither distributional zero nor distributional pole according to Formula (2.5).
  • Figure 2.2 shows the probability density function for $(x\pm 1)^{2}$ according to Formula (2.6), which has a distributional pole at $x=0$. $(0\pm 1)^{2}$ is the $\chi^{2}$ distribution [2]. At the distributional pole, the probability density function resembles a Delta distribution.
@@
In both Figure 2.2 and 2.2, the result probability density function in the $\tilde{z}$ representation resembles the input distribution more when the mode of the distributions is further away from either distributional pole or distributional zero.
@@ 2.3 Statistical Taylor Expansion
Formula (2.5) gives the uncertainty distribution of an analytic function. However, normal scientific and engineering calculations usually do not care about the result distribution, but just few statistics of the result, such as the result mean and deviation [2][3]. Such simple statistics is the result of statistical Taylor expansion. The deviation $\delta x$ in statistical Taylor expansion is also referred to as uncertainty, which may or may not be the measured statistical deviation.
An analytic function $f(x)$ can be precisely calculated in a range by the Taylor series as shown in Formula (2.8). Formula (2.7) defines variance momentum $\zeta(n)$. Using Formula (2.7), Formula (2.9) and (2.3) gives the mean $\overline{f(x)}$ and the variance $\delta^{2}f(x)$ for $f(x)$, respectively. $\overline{f(x)}-f(x)$ is defined as the uncertainty bias which is the effect of the input uncertainty to the result value.
@@
$\delta^{2}f(x)=$ $\overline{(f(x)-\overline{f(x)})^{2}}=\overline{f(x)^{2}}-\overline{f(x)}^{2}$
$=\sum_{n=2}^{\infty}(\delta x)^{n}\sum_{j=1}^{n-1}\frac{f^{(j)}_{x}}{j!}\frac{f^{(n-j)}_{x}}{(n-j)!}\left(\zeta(n)-\zeta(j)\zeta(n-j)\right);$ (2.10)
When $\sigma$ is limited, Formula (2.7) provides statistical bounding $\tilde{x}\in(x-\sigma\delta x,x+\sigma\delta x)$.
Due to uncorrelated uncertainty assumption, Formula (2.12) and (2.3) calculate the mean and variance of Taylor expansion in Formula (2.11), in which $\zeta_{x}(m)$ and $\zeta_{y}(n)$ are the variance momenta for $x$ and $y$, respectively:
@@
$=$ $\sum_{m=0}^{\infty}\sum_{n=0}^{\infty}(\delta x)^{m}(\delta y)^{n}\frac{f^{(m,n)}_{(x,y)}}{m!\;n!}\zeta_{x}(m)\zeta_{y}(n);$
$\delta^{2}f(x,y)=$ $\sum_{m=2}^{\infty}\sum_{n=2}^{\infty}(\delta x)^{m}(\delta y)^{n}\sum_{i=0}^{m}\sum_{j=0}^{n}\frac{f^{(i,j)}_{(x,y)}}{i!\;j!}\frac{f^{(m-i,n-j)}_{(x,y)}}{(m-i)!\;(n-j)!}$
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\left(\zeta_{x}(m)\zeta_{y}(n)-\zeta_{x}(i)\zeta_{x}(m-i)\;\zeta_{y}(j)\zeta_{y}(n-j)\right);$ (2.13)
Although Formula (2.3) is only for 2-dimensional, it can be extended easily to any dimensional.
2.4 Bounding Symmetry
A bounding $[x-\sigma_{-}\delta x,x+\sigma_{+}\delta x]$ is mean-preserving if $\zeta(1)=0$, which has the same mean as the unbounded distribution. Usually, the two positive value $\sigma_{-}$ and $\sigma_{+}$ can be chosen to achieve mean-preserving. With mean-preserving bounding, Formula (2.14) and (2.15) give the result for $x\pm y$, while Formula (2.16) and (2.17) give the result for $xy$:
$\overline{x\pm y}$ $=x\pm y;$ (2.14)
$\delta^{2}(x\pm y)$ $=\zeta_{x}(2)(\delta x)^{2}+\zeta_{y}(2)(\delta y)^{2};$ (2.15)
$\overline{xy}$ $=xy;$ (2.16)
$\delta^{2}(xy)$ $=\zeta_{x}(2)(\delta x)^{2}y^{2}+x^{2}\zeta_{y}(2)(\delta y)^{2}+\zeta_{x}(2)(\delta x)^{2}\zeta_{y}(2)(\delta y)^{2};$ (2.17)
For any input distribution $\rho(\tilde{x},x,\delta x)$ symmetric around its mean $x$, any $[x-\sigma\delta x,x+\sigma\delta x]$ bounding has property $\zeta(2n+1)=0$, which simplifies statistical Taylor expansion further.
@@ 2.5 One Dimensional Examples
$\frac{\overline{e^{x}}}{e^{x}}$ $=1+\sum_{n=1}^{\infty}(\delta x)^{n}\zeta(2n)\frac{1}{n!};$ (2.19)
$\frac{\delta^{2}e^{x}}{(e^{x})^{2}}$ $=\sum_{n=2}^{\infty}(\delta x)^{n}\sum_{j=1}^{n-1}\frac{\zeta(n)-\zeta(j)\zeta(n-j)}{j!\;(n-j)!};$ (2.20)
Formula (2.22) and (2.23) give the mean and variance for $\log(x)$, respectively:
$\log(x+\tilde{x})-\log(x)=\log(1+\frac{\tilde{x}}{x})=\sum_{j=1}^{\infty}\frac{(-1)^{j+1}}{j}\frac{\tilde{x}^{j}}{x^{j}};$ (2.21)
$\overline{\log(x)}$ $=\log(x)+\sum_{n=1}^{+\infty}P(x)^{n}\frac{(-1)^{n+1}\zeta(n)}{n};$ (2.22)
$\delta^{2}\log(x)$ $=\sum_{n=2}^{+\infty}P(x)^{n}\sum_{j=1}^{n-1}\frac{\zeta(n)-\zeta(j)\zeta(n-j)}{j(n-j)};$ (2.23)
@@
$\frac{\overline{x^{c}}}{x^{c}}$ $=1+1\sum_{n=1}^{\infty}P(x)^{n}\zeta(n)\begin{pmatrix}c\\ n\end{pmatrix};$ (2.25)
$\frac{\delta^{2}x^{c}}{(x^{c})^{2}}$ $=\sum_{n=2}^{\infty}P(x)^{n}\sum_{j=1}^{n-1}\begin{pmatrix}c\\ j\end{pmatrix}\begin{pmatrix}c\\ n-j\end{pmatrix}\left(\zeta(n)-\zeta(j)\zeta(n-j)\right);$ (2.26)
$\sin(x+\tilde{x})$ $=\sum_{n=0}^{\infty}\eta(n,x)\frac{\tilde{x}^{n}}{n!};\;\;\;\eta(n,x)\equiv\begin{cases}n=4j:\;\;\;sin(x);\\ n=4j+1:\;\;\;cos(x);\\ n=4j+2:\;\;\;-sin(x);\\ n=4j+3:\;\;\;-cos(x);\end{cases}$ (2.27)
$\overline{\sin(x)}=$ $sinx(x)+\sum_{n=1}^{\infty}(\delta x)^{n}\eta(n,x)\frac{\zeta(n)}{n!};$ (2.28)
$\delta^{2}\sin(x)=$ $\sum_{n=2}^{\infty}(\delta x)^{n}\sum_{j=1}^{n-1}\frac{\eta(j,x)\eta(n-j,x)}{j!(n-j)!}\left(\zeta(n)-\zeta(j)\zeta(n-j)\right);$ (2.29)
The result variance in statistical Taylor expansion shows the nature of the calculation, such as $\delta x\rightarrow P(e^{x})$, $\delta x\rightarrow\delta\sin(x)$, $P(x)\rightarrow P(x^{c})$, and $P(x)\rightarrow\delta\log(x)$.
@@ 2.6 Dependency Tracing
  • Formula (2.30) shows $\delta^{2}(f+g)$, whose dependence tracing can be demonstrated by $\delta^{2}(f-f)=0$, and $\delta^{2}(f(x)+g(y))=\delta^{2}f+\delta^{2}g$, with the latter corresponding to Formula (2.15). Formula (2.34) shows that Formula (2.3) uses Formula (2.30) between any two terms in the Taylor expansion of Formula (2.8). In contrast, Formula (2.35) shows that statistical Taylor expansion can no longer be treated as polynomial due to correlation between $x$ and $x^{n}$.
@@ 2.7 Traditional Execution and Dependency Problem
  • Traditionally, an analytic expression can be broken down into simpler and independent arithmetic operations of negation, addition, multiplication, division, and etc  [^5] Sometimes square root is also regarded as an arithmetic operation.. But such approach causes the dependency problem in both floating-point arithmetic and interval arithmetic. It also causes the dependency problem in statistical Taylor expansion also. For example, if $x^{2}-x$ is calculated as $x^{2}-x$, $x(x-1)$, and $(x-\frac{1}{2})^{2}-\frac{1}{4}$, only $(x-\frac{1}{2})^{2}-\frac{1}{4}$ gives the correct result, while the other two give wrong results for wrong independence assumptions between $x^{2}$ and $x$, or between $x-1$ and $x$, respectively.
  • Traditionally, a large calculation may be broken into sequential steps, such as calculating $f(g(x))$ as $f(y)|_{y=g(x)}$. But this method no longer works in statistical Taylor expansion due to dependency tracing of $g(x)$ inside $f(g(x))$, e.g., $\overline{f(g(x))}\neq\overline{f(y)|_{y=g(x)}}$ and $\delta^{2}f(g(x))\neq\delta^{2}f(y)|_{y=g(x)}$. $f(y)|_{y=g(x)}$ is path-dependent, e.g., $\overline{(\sqrt{x})^{2}}>\overline{\sqrt{x^{2}}}$ and $\delta^{2}(\sqrt{x})^{2}>\delta^{2}\sqrt{x^{2}}$.